hello, I am having some trouble with this question. Can someone please help me.
Miss Harris is a buyer for Plough.She can purchase 100 pounds of chemicals for $97. At this rate ,how much would 2,000 pounds of chemical cost?
I also have another question if that is okay.
10,000euro is equivalent to 0.7615 US dollars. Use a proportion to convert $2,500 to Euro.
I would to thank you for your help in advance.

Found 3 solutions by richwmiller, ssriniv1, stanbon:
Answer by richwmiller(17219)

For various reasons one question at a time. (One reason is it is a website rule.)
There are several ways to deal with the first problem
2000/100=20
97*20=$1940 for 2000 pounds of chemicals
the other is similar
97/100=.97 a pound
2000*.97=$1940

Answer by ssriniv1(1)

100 Pounds of chemical cost $ 97
1 Pound of chemical will cost $ 97/100
2000 Pounds of chemical will cost $ (97/100) * 2000 = $ 1940
btw, is the second question a real-life question or a made up one. 10,000 eur is not equal to 0.7615 USD, it is much more than that. One can still answer this if that is a made question. Do let me know.

Answer by stanbon(75887)

Miss Harris is a buyer for Plough.
She can purchase 100 pounds of chemicals for $97. At this rate ,how much would 2,000 pounds of chemical cost?
-----
cost/weight = cost/weight
---
x/2000 = 97/100
---------
x = 2000(97/100)
---
x = 20*97
---
cost = $1940.00
============================
I also have another question if that is okay.
10,000 euro is equivalent to 0.7615 US dollars. Use a proportion to convert $2,500 to Euro.
----
Note: Your problem is confusing because a Euro is actually
worth more than a US dollar. Today's paper shows the Euro
worth 1.35 dollars. That means the dollar is worth 0.74 euro.
----
Converting 2500 dollars to Euro would be as follows:
Euro/dollars = Euro/dollars
E/2500 = 1/1.35
E = 2500(1/1.35)
# of Euros = 2500/1.35 = 0.74(2500) = 1851
====================================================
Cheers,
Stan H.